The refractive indices of glycerine and diamond with respect to air are 1.4 and 2.4 respectively.calculate the speed of light in glycerine and diamond.From these results, calculate the refractive index of diamond with respect to glycerine. (C=3 into 10^8 m/s).

Grade:9

1 Answers

Sanjana
5 years ago
Refractive index of glycerin=1.4
Refractive index of diamond=2.4
Speed of glycerin=speed of light/refractive index of glycerin=3*10^8/1.4=2.143*10^8
 
Speed of light in diamond=speed of light/refractive index of diamond
=3*10^8/2.4=1.250*10^8
 
Refractive index of diamond w.r.t glycerin=2.4/1.4=1.714